# Explaining Context Length Scaling and Bounds for Language Models

### v1

@@ 1 Introduction
Because of the rapid development of capacity of Language Models and the importance of a long context length in tasks like reasoning, retrieval, etc, past years people have been attempting to extend the context length of Language Models. There have been a variety of methods on supporting long context Language Models(Su et al., 2023; Katharopoulos et al., 2020; Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024). A wide variety of work is proposed to discuss the impact of context length: some shows long irrelevant context would worsen performance for LMs(Xu et al., 2024; Levy et al., 2024); some shows long context would improve performance in a way summarized as Scaling Laws(Xiong et al., 2024); while work in other domain like time series shows long relevant context would hurt performance (Shi et al., 2024). This calls for a more thorough understanding about how context length impact Language Models’ performance..
Previously, theories have been proposed to explain the Scaling Laws with respect to Dataset and Model size(Bahri et al., 2024; Sharma & Kaplan, 2020). However, these theories do not study how context length impact Language Modeling, thus they cannot contribute directly to the problem.
@@
Our theoretical framework can naturally derive phenomena we have observed. For example, our theory derive that, for a certain amount of training data, as context length increases, the neural network first would first behave like the Bayes Model (thus loss decreases); then beyond certain optimal context length, the gap between the trained model and the Bayes Model would increase, hence the validation loss would increase: this is experimentally verified, as shown in Figure 1,
@@
Figure 1: Upper figure: Validation Loss v.s. Context Length, measured on subsets of OpenWebText dataset. We see for each training dataset size, there exists an optimal context length that minimizes pretraining validation loss, which increases with the dataset size (More details can be found in Section 3). Lower figure: similar results obtained on our Synthetic Dataset in Section 4.5. This proves the deduction of our theory.
@@ 2.1 Preliminary: Loss Decomposition and Intrinsic Space
It is common in ML studies to decompose the loss into Bayes Risk (the minimum loss possible, achieved by the theoretically optimal Bayesian Model), and Approximation Loss (the loss measuring the ability of an actually trained model to approximate the Bayesian Model). Specifically for Cross-Entropy loss, we have (please refer to Appendix A for more details):
@@
Bayesian Models are typically related to the Intrinsic Space of data manifold in such a way: the middle-layer or last-layer data representation of a well-trained Bayesian Neural Network is often used to approximate the Intrinsic Space(Cheng et al., 2023). In Section 2.2 we would discuss more how this bridges Bayes Risk with Intrinsic Space, and how context length impacts Bayes Risk.
Approximation Loss, or how well the trained model learns Bayesian Model, is also related to Intrinsic Dimension in the perspective of Scaling Laws. In Section 2.3 we would discuss more about how context length impacts Approximation Loss from this perspective.
We utilize a synthetic dataset to prove concepts in Section 4.
We further derive that, the balance between Bayes Risk and Approximation Loss would lead to an optimal context length which increases with training dataset size. Our theoretical deduction, and experiments on Language and synthetic data are presented in Section 3 and Section 4.5.
@@ 2.2.1 Bayes Risk and Entropy in Intrinsic Space: derived from First Principles
We propose a simple theory model to relate $H(P,P_{l})$ with Intrinsic Dimension $dim(l)$ of intrinsic space $\text{space}_{l}$ of text corpora of length $l$ (for the next-token-prediction task).
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $dim(l_{1})<dim(l_{2})$. This is because longer context contains more information about the next possible token.
@@
  • Assumption 3. Each intrinsic dimension would add $s$ bits of information to the next-token-prediction task, so in all there are $dim(l)*s$ bits of information that can be represented in $\text{space}_{l}$ for next token prediction. Note this does not mean these are the only information in the Intrinsic Space, hence $s$ can be small, or even smaller than $1$.
  • Assumption 4. KL-divergence for Bayes Model of context length $l$, $P_{l}$, with Bayes Model of infinite context length, $P=P_{\infty}$, equals to $s*(dim(\infty)-dim(l))$.
@@
Note by Assumption 1 and 2 we derive that:
@@ 2.2.2 Bayes Risk and Intrinsic Dimension: Experiment Measurement
Figure 2: Bayes Risk v.s. Context Length: Bayes Risk is approximated by Cross Entropy loss measured with LLaMa-3.1 series on OpenWebText, for different context length.
@@
We further use PCA as metric to measure the Intrinsic Dimension of Dataset with respect to context length. Figure 3 provides the relative eigen value degradation in the feature space of LLaMa-3.1-8B, for the last token. We see that larger input length would indeed provide feature with lower degradation in the intrinsic space. Notably, when $15005<idx<$ 1500 the curves is similar to Zip-f distribution ($\log eig=C_{0}-C*\log idx$), and for $4000500<idx<$ 4000 it resembles exponential degradation ($\log eig=C_{0}-C*idx$). Naturally, w ‘Intrinsic Dimension’. However, these formulas are very similar around $idx\approx 1000$, hence cannot provide accurate estimations to the transformation index.
Instead, following previous practice, here we use some threshold to decide the Ine would like to use the transformation index of these two states astrinsic Dimension: $\max_{idx}rela\_eig(idx)\geq\text{threshold}$ is used as the measured Intrinsic Dimensionz. We further evaluate the validity of this method in Point 2 in Section 4.1, on a synthetic Dataset with synthetic data; and explain the linear measurement from an Entropy perspective in Appendix C. Notably, the threshold here is a hyperparameter which is set to constants like $1/20$ in some previous work(Aghajanyan et al., 2020), but we observe here that many thresholds would validate the linear correspondence of Cross Entropy v.s. Intrinsic Dimension, which further enhance the robustness of our result. In detail, as shown in Figure 3 and Figure 4, we use thresholds from $0.002$ to $0.25$ to make our conclusion more robust.
For a certain threshold, we conduct experiments on several context lengths, and measure CE Loss on certain text corpora with these context lengths. We observe quite linear relationship between CE Loss and ID measured (supporting our theory), as shown in Figure 4:
@@
Figure 4: Cross Entropy Loss v.s. Measured Intrinsic Dimension, for LLaMa-3.1-8B on a subset of OpenWebText. Each line represents a certain threshold used to measure ID in the intrinsic space of the used LLM. Different Measurements would give ID values that are linear w.r.t. each other, and they are all linear w.r.t. CE loss.
@@ 3 Deduction: Optimal Context Length and Training Dataset Size
In previous sections we did not specifically discuss the relationship between $A$ and $l$. We consider $l$ where $\partial_{l}L_{CE}=0$ would give us an optimal $l$ with respect to $D$:
@@
We find that $\lim_{D\rightarrow 0}f=-\infty$ and $\lim_{D\rightarrow\infty}f=\infty$. This shows that, for fixed $l$, no matter what $\partial_{l}A$ is, exist some $D$ s.t. $\partial_{l}L_{CE}=0$.
We see that Bayes Risk decreases with $l$, Approximation Loss increases with $l$ but decreases with $D$; the balance between these two losses results in an optimal $l$ that increases with optimal $D$.
We conduct experiments on a subset of OpenWebText with long enough context length. We trained GPT-2 on different context length with different amount of training data, until validation loss increases. We show our results in Figure 1 and Figure 5. Details for our experiment settings are presented in Appendix F.
@@
Figure 5: Openwebtext subset, Validation Loss v.s. Context Length, for different dataset sizes. Different curves represent different amount of training data used. A more readable figure can be found in Figure 1, where the minimum validation loss reachable for each training dataset size is subtracted.
@@ 4.2 Construction of Synthetic Data: the ‘position weighted multitask sparse parity’ dataset
In previous work, a common practice is to mask the leftmost tokens and leave $l$ tokens before the token-to-predict visible to Language Models, as shown in the in Figure 6. Although this may not show the impact of important tokens to final answer perplexity (e.g., it fails to show the importance of the second key info in Figure 6), this method aligns well with our setting of increasing context length.
Figure 6: An example of the ‘two needles in a haystack’ task, similar to those in (Levy et al., 2024). Left: the input to the Language Model, with key information and question visualized in blue. Right: perplexity of the answer token $\langle 8\rangle$ of LLaMa-3.1-8B (horizontal axis) v.s. number of masked leftmost tokens (vertical axis). Although seeing both pieces of information are necessary to answer the question, perplexity rises dramatically only when the first piece of information is masked.
Although the next token to predict might be dependent of several pieces of key information, we see from Figure 6 that the first key token would raise model perplexity.
Inspired by this concept in Figure 6 and the ‘multitask sparse parity dataset’ previously studied in (Michaud et al., 2024; syntheticdatasetorigin), we propose the ‘position weighted multitask sparse parity dataset’. In detail, each input is consisted of $T$ ‘control bits’ and $L$ ‘context bits’, each bit lies in $\{0,1\}$. Each subtask take xor on two certain bits in the context bits, and the answer to some sample is the answer of the only activated subtask, as shown in Figure 7. We set tasks to be of equal probability in the training and testing datasets. To emphasize the importance of context in different positions, for task $t$, we ensure $l(t)=max(context\_bit_{1}(t),context\_bit_{2}(t))$ follows $l(t)\approx 20/(1-t/50)^{1/1.2}$. This ensures,
@@
in which, if we assume $1$ dimension in Intrinsic Space represents $1$ subtask (which has $1$ bit of information about the subtask, and hence $1/T$ bits of information to the output logit), then $s=1/T$, thus $ID(l,1/T)=t(l)$.
4.3 Cross Entropy v.s. Intrinsic Dimension v.s. Context Length
We train a large enough MLP on data generated on the previous tasks, and evaluate our model on the validation dataset. During synthetic data generation we make sure the training dataset and validation dataset does not overlap on any sample. We train until overfitting the training dataset. We assume $1$ dimension in Intrinsic Space can store information about $1$ subtask, hence we take $ID(l)=t(l)$ as its theoretical value here. After training, we obtain such results:
@@
  • ID v.s. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-51.1\pm 1.0$, $C=-1.7*10^{3}\pm 0.3*10^{3}$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE v.s. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-0.015\pm 0.013$, $C=23.8\pm 4.3$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE v.s. ID: $CE\approx g(ID,k,b)$, $k=0.693\pm 1*10^{-5}$, $b\approx 0.0139\pm 4*10^{-7}$, $R^{2}=1-7*10^{-9}$.
As shown, we construct synthetic data such that $ID(l)=ID_{0}-C^{\prime}/l^{\gamma}$, and our measurements show $CE=C-C^{\prime}/l^{\gamma}$. More importantly, for the synthetic data example, Cross Entropy loss is almost perfectly linear with the Intrinsic Dimension as we defined previously. This validates Point 1: Cross Entropy Loss is approximately linear with Intrinsic Dimension; and we have also provided a construction to match the measured $CE(l)\approx C_{0}+C/l^{\gamma}$ relationship.
4.4 Eigen Values in Intrinsic Space v.s. Intrinsic Dimension
In this subsection, we train a model with specialized architecture, allowing us to use the feature representation of a middle layer as feature vector for input context bits, as shown in Figure 8.
@@
From Figure 9 we find that: (1) the neural network would indeed learn the key information in the context bits. For models with different input context length, although their inner dimension are the same ($80$), the representation of inputs in this inner space mainly lies in the first $ID$ dimensions, and eigen values correspond to other dimensions are very small; and (2) there exists such threshold $y^{*}$ that would estimate $ID$ for all context lengths accurately. In detail, to use $y^{*}$ as threshold to estimate $ID$ means that we find $idx\ s.t.\ rela\_eig(idx)>y^{*}$ and $rela\_eig(idx+1)<y^{*}$. We label a rectangle to show the possible $y^{*}$ for each context length, and show that there exists $y^{*}$ to provide accurate estimations in Figure 9.
Figure 9: Illustration of relative eigen value v.s. index, for models trained on different context length. Vertical lines: $x_{l}=ID(l)$ and $x_{l}^{\prime}=ID(l+1)$. For example, a context length $27$ has $16$ subtasks visible, corresponding to $16$ bits in Intrinsic Space. Assuming $1$ dimension in Intrinsic Space represents $1$ bit, the leftmost purple rectangle drawn means a range of $y_{threshold}$ that would provide an accurate estimation of $ID(27)=16$ for context length $27$. Obviously, there exists $y^{*}$ that would provide an estimation of $ID$ for all context lengths, as shown in the figure.
This provides evidence for Point 2 in Section 4.1: we can take some threshold $y^{*}$ to estimate the intrinsic dimension, by obtaining the maximum index of the relative eigen value such that the relative eigen value is larger than $y^{*}$, which would give accurate and consistent estimations.

### v2

@@ 1 Introduction
Because of the rapid development of capacity of Language Models and the importance of a long context length in tasks like reasoning, retrieval, etc, past years people have been attempting to extend the context length of Language Models. There have been a variety of methods on supporting long context Language Models(Su et al., 2023; Katharopoulos et al., 2020; Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024). A wide variety of work is proposed to discuss the impact of context length: some shows long irrelevant context would worsen performance for LMs(Xu et al., 2024; Levy et al., 2024); some shows long context would improve performance in a way summarized as Scaling Laws(Xiong et al., 2024); while work in other domains like time series shows long relevant context would hurt performance (Shi et al., 2024). This calls for a more thorough understanding of how context length affects Language Models’ performance..
Previously, theories have been proposed to explain the Scaling Laws with respect to the data set and the size of the model(Bahri et al., 2024; Sharma & Kaplan, 2020). However, these theories do not study how context length impact Language Modeling, thus they cannot contribute directly to the problem.
@@
Our theoretical framework can naturally derive phenomena that we have observed. For example, our theory derives that, for a certain amount of training data, as the context length increases, the neural network first would first behave like the Bayes Model (thus the loss decreases); then beyond a certain optimal context length, the gap between the trained model and the Bayes Model would increase, hence the validation loss would increase: this is experimentally verified, as shown in Figure 1,
@@
Figure 1: Upper figure: Validation Loss vs. Context Length, measured on subsets of OpenWebText dataset. We see for each training dataset size, there exists an optimal context length that minimizes pretraining validation loss, which increases with the dataset size (More details can be found in Section 3). Lower figure: similar results obtained on our Synthetic Dataset in Section 4.5. This proves the deduction of our theory.
@@ 2.1 Preliminary: Loss Decomposition and Intrinsic Space
It is common in ML studies to decompose the loss into Bayes Risk (the minimum loss possible, achieved by the theoretically optimal Bayesian Model), and Approximation Loss (the loss measuring the ability of a trained model actually to approximate the Bayesian Model). Specifically for Cross-Entropy loss, we have (please refer to Appendix A for more details):
@@
Bayesian Models are typically related to the Intrinsic Space of data manifold in such a way: the middle-layer or last-layer data representation of a well-trained Bayesian Neural Network is often used to approximate the Intrinsic Space(Cheng et al., 2023). In Section 2.2 we discuss more how this bridges Bayes Risk with Intrinsic Space, and how context length impacts Bayes Risk.
Approximation Loss, or how well the trained model learns Bayesian Model, is also related to Intrinsic Dimension in the perspective of Scaling Laws. In Section 2.3 we discuss more about how the context length impacts Approximation Loss from this perspective.
We use a synthetic data set to prove concepts in Section 4.
We further derive that the balance between Bayes Risk and Approximation Loss would lead to an optimal context length which increases with the size of the training dataset. Our theoretical deduction and experiments on language and synthetic data are presented in Section 3 and Section 4.5.
@@ 2.2.1 Bayes Risk and Entropy in Intrinsic Space: derived from First Principles
We propose a simple theory model to relate $H(P,P_{l})$ with the intrinsic dimension $dim(l)$ of the intrinsic space $\text{space}_{l}$ of the text corpora of length $l$ (for the next-token prediction task).
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $dim(l_{1})<dim(l_{2})$. This is because a longer context contains more information about the next possible token.
@@
  • Assumption 3. Each intrinsic dimension would add $s$ bits of information to the next-token prediction task, so there are $dim(l)*s$ bits of information that can be represented in $\text{space}_{l}$ for the next-token prediction. Note this does not mean these are the only information in the Intrinsic Space, hence $s$ can be small, or even smaller than $1$.
  • Assumption 4. KL-divergence for the Bayes ¡+Model of context length $l$, $P_{l}$, with Bayes Model of infinite context length, $P=P_{\infty}$, equal to $s*(dim(\infty)-dim(l))$.
@@
Note that by Assumptions 1 and 2 we derive that:
@@ 2.2.2 Bayes Risk and Intrinsic Dimension: Experiment Measurement
Figure 2: Bayes Risk vs. Context Length: Bayes Risk is approximated by Cross Entropy loss measured with LLaMa-3.1 series on OpenWebText, for different context length.
@@
We further use PCA as a metric to measure the Intrinsic Dimension of Dataset with respect to context length. Figure 3 provides the relative degradation of the eigenvalue in the feature space of LLaMa-3.1-8B, for the last token. We see that larger input length would indeed provide feature with lower degradation in the intrinsic space. Notably, when $15005<idx<$ 1500 the curves is similar to Zip-f distribution ($\log eig=C_{0}-C*\log idx$), and for $4000500<idx<$ 4000 it resembles exponential degradation ($\log eig=C_{0}-C*idx$). Naturally, w ‘Intrinsic Dimension’. However, these formulas are very similar around $idx\approx 1000$, hence cannot provide accurate estimations to the transformation index.
Instead, following previous practice, here we use some threshold to decide the Ine would like to use the transformation index of these two states as Itrinsic Dimension: $\max_{idx}rela\_eig(idx)\geq\text{threshold}$ is used as the measured Intrinsic Dimensionz. We further evaluate the validity of this method in Point 2 in Section 4.1, on a synthetic Dataset with synthetic data; and explain the linear measurement from an Entropy perspective in Appendix C. Notably, the threshold here is a hyperparameter which is set to constants like $1/20$ in some previous work(Aghajanyan et al., 2020), but we observe here that many thresholds would validate the linear correspondence of Cross Entropy vs. Intrinsic Dimension, which further enhance the robustness of our result. In detail, as shown in Figure 3 and Figure 4, we use thresholds from $0.002$ to $0.25$ to make our conclusion more robust.
For a certain threshold, we conduct experiments on several context lengths, and measure CE Loss on certain text corpora with these context lengths. We observe a fairly linear relationship between CE Loss and ID measured (supporting our theory), as shown in Figure 4:
@@
Figure 4: Cross Entropy Loss vs. Measured Intrinsic Dimension, for LLaMa-3.1-8B on a subset of OpenWebText. Each line represents a certain threshold used to measure ID in the intrinsic space of the used LLM. Different Measurements would give ID values that are linear w.r.t. each other, and they are all linear w.r.t. CE loss.
@@ 3 Deduction: Optimal Context Length and Training Dataset Size
In previous sections, we did not specifically discuss the relationship between $A$ and $l$. We consider $l$ where $\partial_{l}L_{CE}=0$ would give us an optimal $l$ with respect to $D$:
@@
We find $\lim_{D\rightarrow 0}f=-\infty$ and $\lim_{D\rightarrow\infty}f=\infty$. This shows that for fixed $l$, no matter what $\partial_{l}A$ is, there exists some $D$ s.t. $\partial_{l}L_{CE}=0$.
We see that Bayes Risk decreases with $l$, Approximation Loss increases with $l$ but decreases with $D$; the balance between these two losses results in an optimal $l$ that increases with the optimal $D$.
We conduct experiments on a subset of OpenWebText with a sufficiently long context length. We trained GPT-2 on different context lengths with different amounts of training data, until the validation loss increases. We show our results in Figure 1 and Figure 5. Details for our experiment settings are presented in the Appendix F.
@@
Figure 5: Openwebtext subset, Validation Loss vs. Context Length, for different dataset sizes. Different curves represent different amount of training data used. A more readable figure can be found in Figure 1, where the minimum validation loss reachable for each training dataset size is subtracted.
@@ 4.2 Construction of Synthetic Data: the ‘position weighted multitask sparse parity’ dataset
In previous work, a common practice is to mask the leftmost tokens and leave $l$ tokens before the token-to-predict visible to Language Models, as shown in Figure 6. Although this may not show the impact of important tokens to final answer perplexity (e.g., it fails to show the importance of the second key info in Figure 6), this method aligns well with our setting of increasing context length.
Figure 6: An example of the ‘two needles in a haystack’ task, similar to those in (Levy et al., 2024). Left: the input to the Language Model, with key information and question visualized in blue. Right: perplexity of the answer token $\langle 8\rangle$ of LLaMa-3.1-8B (horizontal axis) vs. number of masked leftmost tokens (vertical axis). Although seeing both pieces of information are necessary to answer the question, perplexity rises dramatically only when the first piece of information is masked.
Although the next token to predict might depend on several pieces of key information, we see from Figure 6 that the first key token would raise model perplexity.
Inspired by this concept in Figure 6 and the ‘multitask sparse parity dataset previously studied in (Michaud et al., 2024; Barak et al., 2022), we propose the ‘position-weighted multitask sparse parity dataset. In detail, each input consists of $T$ ‘control bits and $L$ ‘context bits, each bit lies in $\{0,1\}$. Each subtask takes xor on two certain bits in the context bits, and the answer to some sample is the answer of the only activated subtask, as shown in Figure 7. We set tasks to have the same probability in the training and testing datasets. To emphasize the importance of context in different positions, for task $t$, we ensure $l(t)=max(context\_bit_{1}(t),context\_bit_{2}(t))$ follows $l(t)\approx 20/(1-t/50)^{1/1.2}$. This ensures,
@@
in which, if we assume that the $1$ dimension in Intrinsic Space represents $1$ subtask (which has $1$ bits of information about the subtask and hence $1/T$ bits of information to the output logit), then $s=1/T$, thus $ID(l,1/T)=t(l)$.
4.3 Cross Entropy vs. Intrinsic Dimension vs. Context Length
We train a large enough MLP on data generated on the previous tasks, and evaluate our model on the validation dataset. During synthetic data generation, we make sure the training dataset and validation dataset do not overlap on any sample. We train until overfitting the training dataset. We assume $1$ dimension in Intrinsic Space can store information about $1$ subtask, hence we take $ID(l)=t(l)$ as its theoretical value here. After training, we obtain such results:
@@
  • ID vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-51.1\pm 1.0$, $C=-1.7*10^{3}\pm 0.3*10^{3}$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-0.015\pm 0.013$, $C=23.8\pm 4.3$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. ID: $CE\approx g(ID,k,b)$, $k=0.693\pm 1*10^{-5}$, $b\approx 0.0139\pm 4*10^{-7}$, $R^{2}=1-7*10^{-9}$.
As shown, we construct synthetic data such that $ID(l)=ID_{0}-C^{\prime}/l^{\gamma}$, and our measurements show $CE=C-C^{\prime}/l^{\gamma}$. More importantly, for the synthetic data example, Cross Entropy loss is almost perfectly linear with the Intrinsic Dimension as we defined previously. This validates Point 1: Cross Entropy Loss is approximately linear with Intrinsic Dimension; and we have also provided a construction to match the measured relationship $CE(l)\approx C_{0}+C/l^{\gamma}$.
4.4 Eigen Values in Intrinsic Space vs. Intrinsic Dimension
In this subsection, we train a model with a specialized architecture, allowing us to use the feature representation of a middle layer as a feature vector for input context bits, as shown in Figure 8.
@@
From Figure 9 we find that: (1) the neural network would indeed learn the key information in the context bits. For models with different input context lengths, although their inner dimensions are the same ($80$), the representation of inputs in this inner space mainly lies in the first $ID$ dimensions, and the eigen values corresponding to other dimensions are very small; and (2) there exists such threshold $y^{*}$ that would estimate $ID$ for all context lengths accurately. In detail, using $y^{*}$ as the threshold to estimate $ID$ means that we find $idx\ s.t.\ rela\_eig(idx)>y^{*}$ and $rela\_eig(idx+1)<y^{*}$. We label a rectangle to show the possible $y^{*}$ for each context length and show that there exists $y^{*}$ to provide accurate estimations in Figure 9.
Figure 9: Illustration of relative eigen value vs. index, for models trained on different context length. Vertical lines: $x_{l}=ID(l)$ and $x_{l}^{\prime}=ID(l+1)$. For example, a context length $27$ has $16$ subtasks visible, corresponding to $16$ bits in Intrinsic Space. Assuming $1$ dimension in Intrinsic Space represents $1$ bit, the leftmost purple rectangle drawn means a range of $y_{threshold}$ that would provide an accurate estimation of $ID(27)=16$ for context length $27$. Obviously, there exists $y^{*}$ that would provide an estimation of $ID$ for all context lengths, as shown in the figure.
This provides evidence for Point 2 in Section 4.1: we can take some threshold $y^{*}$ to estimate the intrinsic dimension, by obtaining the maximum index of the relative eigen value such that the relative eigen value is larger than $y^{*}$, which would give accurate and consistent estimates.